# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 7. Classification theorem
  1. (assump. 5.1) The lattice $L$ is distributive. We define the poset of points (join-irreducibles) to be $J$, so $L={\operatorname{Ideal}_{\textup{f}}}(J)$.
  2. (assump. 5.19) The $J$ is unique-cover-modular.
  3. (assump. 5.22) The weighting $w$ is admissible, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive. The differential degree $r$ is positive.
  4. (assump. 6.1) The lattice cannot be factored into two nontrivial lattices, which is equivalent to assuming that $J$ cannot be partitioned into two disjoint non-empty posets $J=J_{1}\sqcup J_{2}$.
@@
  1. 6.7:
    (para. 2(5)) $J$ is the upward semi-infinite path $B_{i}$ for $i\geq 1$ with $r>0$ and $w(B_{i})=ir$. $L$ is isomorphic to $\mathbb{N}$.
  2. 6.27:
    (para. 2(2)) $J$ is isomorphic to the quadrant of the plane, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers $\geq 0$}\}$, and $L$ is $\mathbb{Y}$, Young’s lattice, the lattice of partitions, with weighting $r$ times the canonical weighting of 1.
  3. 6.29:
@@
  4. 6.31:
    (para. 2(4)) $J$ is isomorphic to the $k$-row strip, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers, $0\leq x$ and $0\leq y<k$}\}$, and $L$ is $\mathbb{Y}_{k}$, the $k$-row Young’s lattice, the lattice of partitions with at most $k$ parts, with weighting $\alpha>0$ times the canonical weighting $w(T_{xy})=k+x-y$.
@@
This is the summation of 6.25, 6.26, 6.28, and 6.30. ∎
@@ 8. Future directions
The strength of the assumptions of our classification theorem (th. 7.1 is disappointing, but they are satisfied by all known Fomin lattices except the one family that is not even distributive, viz., the Young–Fibonacci lattices and cartesian products that include them. In future work, we hope to expand the classification to larger classes. The following considerations come to mind.
There are a number of ways of relaxing the assumption of admissibility.
@@
    Considering sets of values $V$ that are not integers $\mathbb{Z}$ may be of abstract interest but doesn’t seem to be directly relevant to Robinson–Schensted algorithms. However, investigations of “generic” weightings of a lattice can exploit this generality. For instance, looking at weightings of the quadrant as $J$, we could start with the generic formula $w((i,j))=\alpha+\beta r+\gamma i+\delta j$ and see what values of $\alpha$, $\beta$, $\gamma$, and $\delta$ allow the differential criterion to be satisfied. This can be modeled by setting $V=\mathbb{R}^{4}$ and translating $\alpha+\beta r+\gamma i+\delta j$ into $(\alpha,\beta,\gamma,\delta)$ and $r$ into $(0,1,0,0)$. Once the possible weightings for this $V$ are determined, then a homomorphism from $V$ to $\mathbb{Z}$ can be applied and the result weightings that are admissible can be identified. From these admissible weightings Robinson–Schensted algorithms can be instantiated.
@@
    Allowing values of $w$ to be negative (when $V=\mathbb{Z}$) does not seem to be relevant to construction Robinson–Schensted algorithms.
@@
    Allowing values of $w$ to be 0 may be useful as an intermediate step in constructing admissible weightings, but it will reveal no new Fomin lattices since th. 4.16 et al. shows that a Fomin lattice with some edges weighted 0 can be truncated into connected components which are each Fomin lattices with all edges having nonzero weights.
@@
    Allowing $r\leq 0$ (with $V=\mathbb{Z}$) seems to have limited potential as then either $L$ is nontrivial or does not have $\hat{0}$. However, see the discussion of Fomin lattices without $\hat{0}$ in (5).
@@
    Currently, it is impossible to use a Fomin lattice without $\hat{0}$ to build a Fomin lattice with $\hat{0}$, and thus a Robinson–Schensted algorithm. But the truncation theorems (th. 4.16 et al.) suggest it’s worth looking for processes by which they might be exploited. The existence of Fomin lattices without $\hat{0}$ that do not resemble any Fomin lattices with $\hat{0}$ (para. 2(7)(8)) suggests this might be productive.
@@
    Relaxing the assumption that the poset of points is cover-modular seems to be more difficult as it is used in the proof that no element of $L$ has more than two upward or downward covers (lem. 5.27) and a number of the lemmas for global characterization.

### v2

@@ 7. Classification theorem
  1. (assump. 5.1) The lattice $L$ is distributive. We define the poset of points (join-irreducibles) to be $P$, so $L={\operatorname{Ideal}_{\textup{f}}}(P)$.
  2. (assump. 5.13) $P$ is unique-cover-modular.
  3. (assump. 5.24) The weighting $w$ is positive, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive. The differential degree $r$ is positive.
  4. (assump. 6.1) The lattice cannot be factored into two nontrivial lattices, which is equivalent to assuming that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$.
@@
  1. $\mathbb{Y}$:
    (para. 2(2)) $P$ is isomorphic to the quadrant of the plane, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers $\geq 0$}\}$, and $L$ is isomorphic to $\mathbb{Y}$, Young’s lattice, the lattice of partitions, with weighting $r$ times the canonical weighting of 1.
  2. $\mathbb{SY}$:
@@
  3. $\mathbb{Y}_{\bullet}$:
    (para. 2(4)) $P$ is isomorphic to the $k$-row strip for some $k\geq 1$, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers, $0\leq x$ and $0\leq y<k$}\}$, and $L$ is isomorphic to $\mathbb{Y}_{k}$, the $k$-row Young’s lattice, the lattice of partitions with at most $k$ parts, with weighting $r/k$ times the canonical weighting $w(T_{xy})=k+x-y$.
@@ 3. $\mathbb{Y}_{\bullet}$:
Note that conclusion case $\mathbb{Y}_{k}$ comprises case 6.7, the upward semi-infinite path, as $\mathbb{Y}_{1}$ and the case lem. 6.25(1) as $\mathbb{Y}_{k}$ with $k\geq 2$. The theorem is the summation of case 6.7, lem. 6.19, 6.21, 6.22, 6.23, 6.24, and 6.25. ∎
@@ 8. Future directions
The strength of the assumptions of our classification theorem (th. 7.1 is disappointing, but they are satisfied by all known Fomin lattices except the one family that is not even distributive, the Young–Fibonacci lattices and cartesian products that include them. In future work, we hope to expand the classification to larger classes.
There are a number of ways of relaxing the assumption of positivity.
@@
    Considering sets of values $V$ that are not the integers $\mathbb{Z}$ may be of abstract interest but doesn’t seem to be directly relevant to Robinson–Schensted algorithms. However, investigations of “generic” weightings of a lattice can exploit this generality. For instance, looking at weightings of the quadrant as $P$, we could start with the generic formula $w((i,j))=\alpha+\beta r+\gamma i+\delta j$ and see what values of $\alpha$, $\beta$, $\gamma$, and $\delta$ allow the differential criterion to be satisfied. This can be modeled by setting $V=\mathbb{R}^{4}$ and translating $\alpha+\beta r+\gamma i+\delta j$ into $(\alpha,\beta,\gamma,\delta)$ and $r$ into $(0,1,0,0)$. Once the possible weightings for this $V$ are determined, then homomorphisms from $V$ to $\mathbb{Z}$ can be considered and the resulting weightings that are positive can be identified. From these positive weightings Robinson–Schensted algorithms can be constructed.
@@
    Allowing values of $w$ to be negative (when $V=\mathbb{Z}$) does not seem to be relevant to constructing Robinson–Schensted algorithms.
@@
    Allowing values of $w$ to be 0 may be useful as an intermediate step in constructing positive weightings, but it will reveal no new Fomin lattices since th. 4.16 et al. show that a Fomin lattice with some edges weighted 0 can be truncated into connected components which are each Fomin lattices with all edges having nonzero weights.
@@
    Allowing $r\leq 0$ (with $V=\mathbb{Z}$) seems to have limited potential as then either $L$ is trivial or does not have $\hat{0}$. However, see the discussion of Fomin lattices without $\hat{0}$ in (5).
@@
    Currently, there is no known way to use a Fomin lattice without $\hat{0}$ to build a Fomin lattice with $\hat{0}$, and thus a Robinson–Schensted algorithm. But the truncation theorems (th. 4.16 et al.) suggest it’s worth looking for processes by which they might be exploited.
  2. (6)
    There are a considerable number of Fomin lattices which have all values of $w$ positive but $r=0$ (and necessarily, no $\hat{0}$): See para. 2(6)(8)(10)(11). There is no obvious application of these to Robinson–Schensted algorithms, as any cartesian product involving any of them has no $\hat{0}$ and cannot be used in Robinson–Schensted algorithms. But this class of lattices may have a different combinatorial significance.
@@
    Relaxing the assumption that the poset of points is cover-modular seems to be more difficult as it is used in the proof that no element of $L$ has more than two upward or downward covers (lem. 5.30) and a number of the lemmas for global characterization.
@@
    On the other hand, the technique used in [Stan2012a]*Exer. 3.51(a) soln. can be restated to provide a way to analyze an arbitrary differential distributive lattice, or conversely, to provide a nondeterministic algorithm for constructing all differential distributive lattices for a wide range of definitions of “differential”. This can be used to prove that no point is covered by three distinct points (lem. 5.30) and that there is a unique minimum point (lem. 6.3). It may be possible to use this technique to prove additional lemmas used in this analysis without assuming unique-cover-modularity.
@@
  2. (11)
    Richard Stanley asks how much of this analysis can be applied to sequentially differential posets[Stan1990a]*sec. 2141414Sequentially differential posets have a different differential degree for each grade of the poset. that are distributive lattices.
  3. (12)
    Jonathan Farley[Far2003a]*sec. 11 classifies all finitary distributive lattices which have a cover function, where there is a function $f:\mathbb{N}\rightarrow\mathbb{N}$ and each element of the lattice that covers $n$ elements is covered by $f(n)$ elements. This suggests extending his analysis can be extended to the weighted case, where
[…]